The length of a rectangle is 6 feet more than twice the width. If the length is 24 feet, what is the width?
Answer by mananth(16946)

let width be x feet
length = 2x+6
...
but length = 24
..
2x+6 =24
-6
2x+6-6 = 24-6
2x=18
/2
x=9 the width
m.ananth@hotmail.ca